Consider the reaction represented with the following equation Na2Co3+2Hcl ———> 2Nacl +H2O+CO2. What volume of 0.02 moledm-3 Na2(03Ca0) would be require to completely neutraize 40cm3 of 0.10 moldm-3?


1
Expert's answer
2019-12-10T07:44:29-0500
"Na_2CO_3+2HCl = 2NaCl + H_2O + CO2"

"n=c\times V"

"n(HCl) = 0.10\times 0.040 = 0.004 mol"

According to equation "n(Na_2CO_3):n(HCl) = 1:2" , then "n(Na_2CO_3) =\frac{n(HCl)}{2} = \frac{0.004}{2}= 0.002 mol"


"V(Na_2CO_3) = \frac{n}{c} = \frac{0.002}{0.02} = 0.1 L = 100 ml"